# Unbiased Regression-Adjusted Estimation of Average Treatment Effects in Randomized Controlled Trials

### v1

@@ B.1 Omitted Proofs of Section˜3
$\left\|\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda}\right\|_{2}\leq\left\|\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}\right\|_{2}+\left\|\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda}\right\|_{2}\xrightarrow{p}0,$
which completes the proof. ∎
@@
By Lemma˜3, for all $i\in\mathbb{N}$, $\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\xrightarrow{p}\boldsymbol{\beta}^{*}_{\lambda}$. Hence, by Slutsky’s theorem,
${\widehat{\tau}}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}\frac{d_{i}\big(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\big)}{p_{i}}-\frac{1}{n}\sum_{i=1}^{n}\frac{(1-d_{i})\big(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}\big)}{1-p_{i}}\;\xrightarrow{d}\;{\widehat{\tau}}_{\textup{LHT}}^{*},$
where
${\widehat{\tau}}_{\textup{LHT}}^{*}:=\frac{1}{n}\sum_{i=1}^{n}\frac{d_{i}\big(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{*}_{\lambda}\big)}{p_{i}}-\frac{1}{n}\sum_{i=1}^{n}\frac{(1-d_{i})\big(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}^{*}_{\lambda}\big)}{1-p_{i}}.$
Using independence of $d_{i}$’s under simple random assignment together with ${\mathbb{E}}[d_{i}-p_{i}]=0$, we obtain
@@
${\rm Var}\big({\widehat{\tau}}_{\textup{LHT}}^{*}\big)=\frac{1}{n^{2}}\,\big\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}^{*}-{\boldsymbol{\mu}}\big\|_{2}^{2}.$
By Assumptions˜1 and 3, the summands of ${\widehat{\tau}}_{\textup{LHT}}^{*}$ are independent, mean-zero after centering at $\tau$, and uniformly bounded in absolute value. Hence the Lindeberg condition holds, and by the Lindeberg–Feller CLT (e.g., Theorems 5.2 and 5.10 of white),
$\sqrt{n}\,\big({\widehat{\tau}}_{\textup{LHT}}^{*}-\tau\big)\;\xrightarrow{d}\;\mathcal{N}\!\left(0,\;\lim_{n\to\infty}\frac{1}{n}\,\big\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}^{*}-{\boldsymbol{\mu}}\big\|_{2}^{2}\right).$
It remains to show
$\sqrt{n}\,\big({\widehat{\tau}}_{\textup{LHT}}-{\widehat{\tau}}_{\textup{LHT}}^{*}\big)\;\xrightarrow{p}\;0,$
so that the same limit law transfers to ${\widehat{\tau}}_{\textup{LHT}}$. Define
$a_{i}:=\frac{d_{i}}{p_{i}}-\frac{1-d_{i}}{1-p_{i}}=\frac{d_{i}-p_{i}}{p_{i}(1-p_{i})},\qquad\Delta_{i}:=\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda}.$
Then
${\widehat{\tau}}_{\textup{LHT}}-{\widehat{\tau}}_{\textup{LHT}}^{*}=-\,\frac{1}{n}\sum_{i=1}^{n}a_{i}\,{\bm{x}}_{i}^{\top}\Delta_{i}.$
Split $\Delta_{i}=(\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda})+(\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda})$. For the first part, by the leave-one-out identity (cf. (23)) and the bounds already established in the previous lemma,
$\max_{i\leq n}\big\|\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}\big\|_{2}=O_{p}\!\big(n^{-3/4}\big).$
Since $\|a_{i}\|_{\infty}\leq m^{-1}(1-m)^{-1}$ and $\|{\bm{x}}_{i}\|_{2}=O(1)$ by Assumptions˜1 and 3, we get
@@
$\sqrt{n}\left|\frac{1}{n}\sum_{i=1}^{n}a_{i}\,{\bm{x}}_{i}^{\top}\big(\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}\big)\right|\leq\sqrt{n}\,\frac{1}{n}\sum_{i=1}^{n}O(1)\cdot O_{p}\!\big(n^{-3/4}\big)=O_{p}\!\big(n^{-1/4}\big)\;\to\;0.$
For the second part,
@@
$\sqrt{n}\left|\frac{1}{n}\sum_{i=1}^{n}a_{i}\,{\bm{x}}_{i}^{\top}\big(\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda}\big)\right|=\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}a_{i}\,{\bm{x}}_{i}\right\|_{2}\cdot\big\|\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda}\big\|_{2}.$
By a standard CLT for independent summands with bounded second moments,
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}a_{i}{\bm{x}}_{i}=O_{p}(1).$
Moreover, under Assumptions˜1 and 2, the ridge estimator with fixed $\lambda\geq 0$ satisfies $\|\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}^{*}_{\lambda}\|_{2}=O_{p}(n^{-1/2})$. Hence the product is $O_{p}(n^{-1/2})\to 0$. Combining the two parts yields $\sqrt{n}\,({\widehat{\tau}}_{\textup{LHT}}-{\widehat{\tau}}_{\textup{LHT}}^{*})\xrightarrow{p}0$. Therefore, by the asymptotic equivalence lemma, e.g., (rao73, pg. 122), the limit distribution for $\sqrt{n}({\widehat{\tau}}_{\textup{LHT}}-\tau)$ coincides with that of $\sqrt{n}({\widehat{\tau}}_{\textup{LHT}}^{*}-\tau)$, namely a mean-zero Gaussian with variance $\lim_{n\to\infty}\frac{1}{n}\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}^{*}-{\boldsymbol{\mu}}\|_{2}^{2}$. ∎

### v2

@@ B.1 Omitted Proofs of Section˜3
$A_{n}$ $=\frac{\sqrt{n}(\widehat{\tau}_{\textup{LHT}}^{*}-\tau)}{\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}/\sqrt{n}},~~B_{n}=\frac{\sqrt{n}(\widehat{\tau}_{\textup{LHT}}-\widehat{\tau}_{\textup{LHT}}^{*})}{\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}/\sqrt{n}}.$
Step 1: Oracle CLT.
Under independent treatment assignment with probabilities bounded away from zero and one, bounded outcomes, and uniformly bounded covariates, the random variables $\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})$ are bounded. Therefore
${\rm Var}\!\left(\sqrt{n}(\widehat{\tau}_{\textup{LHT}}^{*}-\tau)\right)=\frac{1}{n}\sum_{i=1}^{n}{\rm Var}(\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}))=\frac{1}{n}\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}^{2},$
by definition of $\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}$. Moreover, the assumptions imply that the random variables $\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})-\tau_{i}$ are bounded. Let $C_{1}$ be a constant such that for all $i\in\mathbb{N}$, $\left|\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})-\tau_{i}\right|\leq C_{1}$. Then for any constant $\delta>0$,
$\frac{1}{{\rm Var}\!\left(n(\widehat{\tau}_{\textup{LHT}}^{*}-\tau)\right)^{1+\delta/2}}\sum_{i=1}^{n}{\mathbb{E}}\left[\left|\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})-\tau_{i}\right|^{2+\delta}\right]$
$\leq\frac{C_{1}^{\delta}}{{\rm Var}\!\left(n(\widehat{\tau}_{\textup{LHT}}^{*}-\tau)\right)^{1+\delta/2}}\sum_{i=1}^{n}{\mathbb{E}}\left[\left|\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})-\tau_{i}\right|^{2}\right]$
$=\frac{C_{1}^{\delta}}{{\rm Var}\!\left(n(\widehat{\tau}_{\textup{LHT}}^{*}-\tau)\right)^{\delta/2}}=\frac{C_{1}^{\delta}}{\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}^{\delta}}.$
Thus since $\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}\to\infty$ and $C_{1}$ is a constant, we have
$\lim_{n\to\infty}\frac{1}{{\rm Var}\!\left(n(\widehat{\tau}_{\textup{LHT}}^{*}-\tau)\right)^{1+\delta/2}}\sum_{i=1}^{n}{\mathbb{E}}\left[\left|\frac{z_{i}}{q_{i}}(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})-\tau_{i}\right|^{2+\delta}\right]=0.$
Therefore, by the Lyapunov central limit theorem (Billingsley (2012)),
$A_{n}=\frac{\sqrt{n}(\widehat{\tau}_{\textup{LHT}}^{*}-\tau)}{\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}/\sqrt{n}}\xrightarrow{d}\mathcal{N}(0,1).$
Step 2: Negligibility of the plug-in error.
Then
$\widehat{\tau}_{\textup{LHT}}^{*}-\widehat{\tau}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\,{\bm{x}}_{i}^{\top}(\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\boldsymbol{\beta}_{\lambda}).$
Therefore by triangle inequality and Cauchy–Schwarz inequality, we have
$\left|\widehat{\tau}_{\textup{LHT}}^{*}-\widehat{\tau}_{\textup{LHT}}\right|$ $\leq\frac{1}{n}\left|\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\,{\bm{x}}_{i}^{\top}(\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}_{\lambda})\right|+\frac{1}{n}\left|\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\,{\bm{x}}_{i}^{\top}(\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda})\right|$
$=\frac{1}{n}\left|({\bm{X}}^{\top}({\bm{z}}/\boldsymbol{q}))^{\top}(\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}_{\lambda})\right|+\frac{1}{n}\left|\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\,{\bm{x}}_{i}^{\top}(\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda})\right|$
$\leq\frac{1}{n}\left\|({\bm{X}}^{\top}({\bm{z}}/\boldsymbol{q}))\right\|_{2}\left\|\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}_{\lambda}\right\|_{2}+\frac{1}{n}\left|\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\,{\bm{x}}_{i}^{\top}(\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda})\right|.$
We have
$\left\|\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}_{\lambda}\right\|_{2}$ $=\left\|(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{X}}}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})\right\|_{2}=O(n^{-1})\left\|\widetilde{{\bm{X}}}^{\top}({\bm{z}}{\bm{t}}/\boldsymbol{q})\right\|_{2}$
$=O_{p}(k^{1/2}n^{-1/2}),$ (17)
where the last line holds using Assumptions˜1 and 2 and Lemma˜8. Moreover, invoking Lemma˜8 also gives $\left\|({\bm{X}}^{\top}({\bm{z}}/\boldsymbol{q}))\right\|_{2}=O_{p}(\sqrt{n})$, concluding $\frac{1}{n}\left\|({\bm{X}}^{\top}({\bm{z}}/\boldsymbol{q}))\right\|_{2}\left\|\widehat{\boldsymbol{\beta}}_{\lambda}-\boldsymbol{\beta}_{\lambda}\right\|_{2}=O_{p}(1/n)$. Moreover,
$\widehat{\boldsymbol{\beta}}_{\lambda}-\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}=\frac{(\widetilde{{\bm{X}}}^{\top}\widetilde{{\bm{X}}}+\lambda{\bm{I}})^{-1}\widetilde{{\bm{x}}}_{i}(\widetilde{y}_{i}-\widetilde{{\bm{x}}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda})}{1-\widetilde{h}_{\lambda ii}}=O(n^{-1}),$
Therefore
$\frac{1}{n}\left|\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\,{\bm{x}}_{i}^{\top}(\widehat{\boldsymbol{\beta}}^{(-i)}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda})\right|=O(n^{-1})$
Combining the above, we have
$\left|\widehat{\tau}_{\textup{LHT}}^{*}-\widehat{\tau}_{\textup{LHT}}\right|=O_{p}(n^{-1}).$
Step 3: Conclusion.
Since $A_{n}\xrightarrow{d}\mathcal{N}(0,1)$ and $B_{n}\xrightarrow{p}0$, Slutsky’s Theorem implies
$\frac{\sqrt{n}(\widehat{\tau}_{\textup{LHT}}-\tau)}{\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}/\sqrt{n}}\xrightarrow{d}\mathcal{N}(0,1).$
Proof of Lemma˜2.
To observe consistency, note that
$\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\bigl(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}\bigr)=\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\bigl(y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\gamma}_{n}\bigr)+\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}{\bm{x}}_{i}^{\top}\bigl(\boldsymbol{\gamma}_{n}-\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}\bigr).$
Since $\boldsymbol{\gamma}_{n}$ is deterministic, the first term on the right-hand side above is an unbiased estimate of $\tau$. Moreover, for the second term, for some constant $C$, we have
$\left|\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}{\bm{x}}_{i}^{\top}\bigl(\boldsymbol{\gamma}_{n}-\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}\bigr)\right|$ $\leq\frac{1}{n}\sum_{i=1}^{n}\left|\frac{z_{i}}{q_{i}}\right|\left\|{\bm{x}}_{i}\right\|_{2}\left\|\boldsymbol{\gamma}_{n}-\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}\right\|_{2}$
$\leq C\max_{i\in[n]}\left\|\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}-\boldsymbol{\gamma}_{n}\right\|_{2},$ (18)
where the second inequality follows from Assumption˜1. Therefore, by (6),
$\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}{\bm{x}}_{i}^{\top}\bigl(\boldsymbol{\gamma}_{n}-\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}\bigr)\xrightarrow{p}0,$
and ${\widehat{\tau}}_{\textup{HLA}}$ is consistent. For the variance, note that
${\widehat{\tau}}_{\textup{HLA}}-\tau=\frac{1}{n}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\bigl(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}\bigr).$
Therefore
$\sqrt{n}({\widehat{\tau}}_{\textup{HLA}}-\tau)$ $=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\left(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\gamma}_{n}\right)+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\left({\bm{x}}_{i}^{\top}\boldsymbol{\gamma}_{n}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\gamma}}_{n}\right)$ (19)
$+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\left({\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\gamma}}_{n}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}\right).$
By Cauchy–Schwarz inequality,
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\bigl({\bm{x}}_{i}^{\top}\boldsymbol{\gamma}_{n}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\gamma}}_{n}\bigr)$ $=\frac{1}{\sqrt{n}}\left(\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}{\bm{x}}_{i}\right)^{\top}(\boldsymbol{\gamma}_{n}-\widehat{\boldsymbol{\gamma}}_{n})$
$\leq\frac{1}{\sqrt{n}}\left\|\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}{\bm{x}}_{i}\right\|_{2}\left\|\boldsymbol{\gamma}_{n}-\widehat{\boldsymbol{\gamma}}_{n}\right\|_{2}.$
Note that by Assumptions˜1 and 2 and Lemma˜8, $\left\|\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}{\bm{x}}_{i}\right\|_{2}=O_{p}(\sqrt{n})$. Therefore since $\left\|\widetilde{{\bm{X}}}\boldsymbol{\gamma}_{n}-{\boldsymbol{\mu}}\right\|_{2}=\Omega(\sqrt{n})$ by (6),
$\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\bigl({\bm{x}}_{i}^{\top}\boldsymbol{\gamma}_{n}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\gamma}}_{n}\bigr)}{\left\|\widetilde{{\bm{X}}}\boldsymbol{\gamma}_{n}-{\boldsymbol{\mu}}\right\|_{2}/\sqrt{n}}\xrightarrow{p}0.$ (20)
Moreover by triangle inequality and Cauchy–Schwarz inequality,
$\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\bigl({\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\gamma}}_{n}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}\bigr)\right|$ $\leq\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\left\|\frac{z_{i}}{q_{i}}{\bm{x}}_{i}\right\|_{2}\left\|\widehat{\boldsymbol{\gamma}}_{n}-\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}\right\|_{2}$
$\leq\frac{1}{\sqrt{n}}\left(\sum_{i=1}^{n}\left\|\frac{z_{i}}{q_{i}}{\bm{x}}_{i}\right\|_{2}\right)\max_{i\in[n]}\left\|\widehat{\boldsymbol{\gamma}}_{n}-\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}\right\|_{2}$
$=o_{p}(1),$
where the last equality follows from Assumptions˜1 and 2 and (6). Therefore since $\left\|\widetilde{{\bm{X}}}\boldsymbol{\gamma}_{n}-{\boldsymbol{\mu}}\right\|_{2}=\Omega(\sqrt{n})$,
$\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\bigl({\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\gamma}}_{n}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\gamma}}_{n}^{(-i)}\bigr)}{\left\|\widetilde{{\bm{X}}}\boldsymbol{\gamma}_{n}-{\boldsymbol{\mu}}\right\|_{2}/\sqrt{n}}\xrightarrow{p}0.$ (21)
Now similar to the proof of Theorem˜2, the Lyapunov central limit theorem (Billingsley (2012)) implies that
$\frac{\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{z_{i}}{q_{i}}\bigl(r_{i}\mu_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\gamma}_{n}\bigr)}{\left\|\widetilde{{\bm{X}}}\boldsymbol{\gamma}_{n}-{\boldsymbol{\mu}}\right\|_{2}/\sqrt{n}}\xrightarrow{d}\mathcal{N}(0,1).$ (22)
Therefore by (19), (20), (21), and (22), and Slutsky’s theorem,
$\frac{\sqrt{n}({\widehat{\tau}}_{\textup{HLA}}-\tau)}{\left\|\widetilde{{\bm{X}}}\boldsymbol{\gamma}_{n}-{\boldsymbol{\mu}}\right\|_{2}/\sqrt{n}}\xrightarrow{d}\mathcal{N}(0,1).$
Proof of Theorem˜3.
Note that $\boldsymbol{\beta}_{0}=\mathop{\mathrm{argmin}\,}\limits_{\boldsymbol{\gamma}}\left\|\widetilde{{\bm{X}}}\boldsymbol{\gamma}-{\boldsymbol{\mu}}\right\|_{2}$ where $\boldsymbol{\beta}_{0}$ is $\boldsymbol{\beta}_{\lambda}$ for $\lambda=0$. Therefore the sequence $\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{0}-{\boldsymbol{\mu}}\right\|_{2}$ is pointwise smaller than or equal to the sequence $\left\|\widetilde{{\bm{X}}}\boldsymbol{\gamma}_{n}-{\boldsymbol{\mu}}\right\|_{2}$ for any choice of $\boldsymbol{\gamma}_{1},\boldsymbol{\gamma}_{2},\ldots$. Therefore the results is implied by Lemmas˜2 and 2. ∎
@@
Let $s_{i}=z_{i}\left(\frac{y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{q_{i}}\right)-\,\tau$,
$\widehat{V}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}\widehat{s}_{i}^{2},~~\text{and}~~\widetilde{V}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}s_{i}^{2}.$
Let $\tau_{i}=y^{(1)}_{i}-y^{(0)}_{i}$. Then
@@
${\mathbb{E}}\left[\widetilde{V}_{\textup{LHT}}\right]$ $=\frac{1}{n}\sum_{i=1}^{n}{\mathbb{E}}\left[\left(z_{i}\frac{y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{q_{i}}-\tau\right)^{2}\right]$
$=\frac{1}{n}\sum_{i=1}^{n}{\mathbb{E}}\left[\left(z_{i}\frac{y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{q_{i}}-\tau_{i}\right)^{2}\right]$
$+2{\mathbb{E}}\left[z_{i}\frac{y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{q_{i}}-\tau_{i}\right](\tau_{i}-\tau)+(\tau_{i}-\tau)^{2}.$
Therefore since
@@
${\mathbb{E}}\left[\left(z_{i}\frac{y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{q_{i}}-\tau_{i}\right)^{2}\right]=(\mu_{i}-\widetilde{{\bm{x}}}_{i}^{\top}\boldsymbol{\beta}_{\lambda})^{2},~~\text{and}~~{\mathbb{E}}\left[z_{i}\frac{y_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{q_{i}}-\tau_{i}\right]=0,$
we have
@@
${\mathbb{E}}\left[\widetilde{V}_{\textup{LHT}}\right]=\frac{1}{n}\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}^{2}+\frac{1}{n}\sum_{i=1}^{n}(\tau_{i}-\tau)^{2}\geq\frac{1}{n}\left\|\widetilde{{\bm{X}}}\boldsymbol{\beta}_{\lambda}-{\boldsymbol{\mu}}\right\|_{2}^{2}=n{\mathbb{E}}\left[(\widehat{\tau}_{\textup{LHT}}^{*}-\tau)^{2}\right].$
Therefore $\widetilde{V}_{\textup{LHT}}$ is a conservative estimate for the variance of $\sqrt{n}\widehat{\tau}_{\textup{LHT}}^{*}$. Also note that we have ${\rm Var}(\widetilde{V}_{\textup{LHT}})=\frac{1}{n^{2}}\sum_{i=1}^{n}{\rm Var}(s_{i}^{2})=O(n^{-1})$. So Using Chebyshev we get
$\widetilde{V}_{\textup{LHT}}-{\mathbb{E}}[\widetilde{V}_{\textup{LHT}}]\xrightarrow{p}0.$
For $\widehat{V}_{\textup{LHT}}$ we can write:
$\widehat{V}_{\textup{LHT}}-\widetilde{V}_{\textup{LHT}}=\frac{1}{n}\sum_{i=1}^{n}\Delta_{i}(\widehat{s}_{i}+s_{i}),$
where $\Delta_{i}=\widehat{s}_{i}-s_{i}=\frac{z_{i}}{q_{i}}{\bm{x}}_{i}^{\top}(\boldsymbol{\beta}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})-({\widehat{\tau}}_{\textup{LHT}}-\tau)$. Using the proof of Theorem˜2, it follows that $\left\|\boldsymbol{\beta}_{\lambda}-\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\right\|_{2}=O_{p}(n^{-1/2})$, and ${\widehat{\tau}}_{\textup{LHT}}-\tau=O_{p}(n^{-1/2})$. Also we have $s_{i}=O(1)$. These two results imply that $\widehat{s}_{i}=O_{p}(1)$. These in conclusoin give $\widehat{V}_{\textup{LHT}}-\widetilde{V}_{\textup{LHT}}\xrightarrow{p}0$, and therefore we get $\widehat{V}_{\textup{LHT}}-{\mathbb{E}}[\widetilde{V}_{\textup{LHT}}]\xrightarrow{p}0$. Finally, since $V_{n}={\mathbb{E}}[\widetilde{V}_{\textup{LHT}}]$ is bounded away from zero for all $n\in\mathbb{N}$, $\widehat{V}_{\textup{LHT}}-{\mathbb{E}}[\widetilde{V}_{\textup{LHT}}]\xrightarrow{p}0$ implies
$\frac{\widehat{V}_{\textup{LHT}}}{V_{n}}\xrightarrow{p}1.$